# Distribution of merging and post-merging galaxies in nearby galaxy clusters

### v1

@@ Abstract
We study the incidence and spatial distribution of galaxies that are currently undergoing gravitational merging (M) or that have signs of a post merger (PM) in six galaxy clusters (A754, A2399, A2670, A3558, A3562, and A3716) within the redshift range, 0.05 0.08. To this aim, we obtained Dark Energy Camera (DECam) mosaics in , , and -bands covering up to of the clusters, reaching 28 mag/arcsec surface brightness limits. We visually inspect color-composite images of volume-limited () cluster-member galaxies to identify whether galaxies are of M or PM types. We find 4% M-type and 7% PM-type galaxies in the galaxy clusters studied. By adding spectroscopic data and studying the projected phase space diagram (PPSD) of the projected clustocentric radius and the line-of-sight velocity, we find that PM-type galaxies are more virialized than M-type galaxies, having 1–5% point higher fraction within the escape-velocity region, while the fraction of M-type was 10% point higher than PM-type in the intermediate environment. Similarly, on a substructure analysis, M types were found in the outskirt groups, while PM types populated groups in ubiquitous regions of the PPSD. Adopting literature-derived dynamical state indicator values, we observed a higher abundance of M types in dynamically relaxed clusters. This finding suggests that galaxies displaying post-merging features within clusters likely merged in low-velocity environments, including cluster outskirts and dynamically relaxed clusters.
@@ 1 Introduction
The importance of mergers in driving the evolution of galaxies within clusters, where effects such as ram-pressure stripping (Gunn & Gott, 1972) and tidal interactions (Jones et al., 2000; Gnedin, 2003) begin to play an increasingly important role, is a matter of discussion. Although clusters represent particularly dense environments, the incidence of mergers among their members is not expected to be high due to high relative velocities. Analytically, the energy input and dynamical friction of a collision scale with the inverse square of the relative velocity of the interacting galaxies (Binney & Tremaine, 1987), such that fast encounters, which become preponderant in denser cluster environments, are less likely to result in a merger than slow encounters, which are more common in the field (Makino & Hut, 1997; Mihos, 2003). Studies such as Sheen et al. (2012) (S12 hereafter) and Oh et al. (2018) (O18 hereafter) report, indeed, that only around four percent of the early-type galaxies (ETGs) found in cluster environments appear to be actively undergoing major mergers. Oh et al. (2018) report only a marginal change in the fraction of ongoing mergers, which increases from 4% to 7.7% when going from 2 to 2 in clustocentric distance, concluding that merger processes are unlikely to significantly impact galaxy evolution within the virialized region of a cluster. McIntosh et al. (2008) found also negative correlation of the major merging frequency of massive satellite galaxies with the host halo mass. Their sample of SDSS satellite galaxies with a stellar mass M dropped their major merging frequency from 3% in halos with their mass around to almost 0% in cluster-scale halos with .
@@
Given the apparently negligible merger rates in cluster environments, the high proportion of cluster galaxies that exhibit post-merging features is therefore unexpected. Deep optical surveys of galaxies in dense cluster environments show a high fraction of post-merging features (25% of galaxies brighter than 20 in the band in the red sequences in S12; 20% of galaxies with similar brightnesses in O18), with abundances comparable to that of the field (see e.g. van Dokkum, 2005; Kado-Fong et al., 2018; Bílek et al., 2020; Sola et al., 2022; Trujillo et al., 2021; Martin et al., 2022; Valenzuela & Remus, 2022). S12 attribute this result to migration of galaxy populations that have undergone mergers prior to entering the cluster (pre-processing). The work of Yi et al. (2013) and Ji et al. (2014) lend support to this argument, showing that, merger features continue to be detectable for about 4 Gyr at a limiting surface brightness of 28 mag arcsec in their hydrodynamic simulations of major mergers between Sa- and Sb-type galaxies. Evidence of pre-processing is also seen at higher redshifts. For instance, van Dokkum et al. (1999) find that merging galaxies are found preferentially in the outskirts of galaxy clusters at , and Olave-Rojas et al. (2018) find high quenching efficiencies in galaxies at large distances from the cluster centers at .
@@
Besides, clusters are consistently accreting groups. Das et al. (2023) found a higher level of star formation (SF) and bluer colors in low-mass SDSS galaxy pairs in filaments and sheets around galaxy clusters. Kleiner et al. (2014) found more asymmetric galaxies in the inner region of A1664 where a merging group most likely passed the pericenter, and they argued that a merger between halos of a group and a cluster enhanced galaxy-galaxy interactions (see also Vijayaraghavan & Ricker, 2013). Mihos (2003) also stressed that the infalling galaxy groups through the “cosmic web” would allow galaxies to interact with each other strongly by slow encounters. Lokas (2023) also found galaxy-galaxy interactions in a galaxy group prior to cluster infall in the IllustridTNG100 simulation. Analyses of substructures in/outside of clusters seem to be necessary to study galaxy-galaxy interactions in merging clusters (i.e., Olave-Rojas et al., 2023).
@@
In this paper, we study the abundances of gravitationally interacting or post-merger galaxies using deep (28 mag arcsec)5 and wide (3 deg; 3) photometric data sets (in , , and ) of six Abell clusters: A754, A2399, A2670, A3558, A3562, and A3716 in a redshift range, 0.05 0.08 collected using the Dark Energy Camera (DECam; DePoy et al., 2008). In combination with public spectroscopic data, we study the prevalence of these features in substructures and in regions of the projected phase-space diagram. This is done to understand the evolution of galaxies over the course of their accretion onto the clusters.
@@ 2.1 Photometric measurements
Optical mosaics in , , and bands of the field of views (FoVs) around seven Abell galaxy clusters were taken using the Dark Energy Camera (DECam; e.g., Flaugher et al., 2015) installed on the Blanco 4m Telescope at CTIO from April 10th to 11th in 2013 (Proposal ID: 2013A-0612, PI: Y. Sheen) and August 19th to 22nd in 2014 (Proposal ID: 2014B-0608, PI: Y. Jaffé), as presented in Table 1 (for A3574 and A3659, visit the website111https://data.kasi.re.kr/vo/DECam_catalogs/). DECam features a 3 deg FoV consisting of sixty-two 2k 4k pixel CCDs, with a total of 520 million pixels for each FITS222Flexible Image Transport System (Wells et al., 1981; Hanisch et al., 2001) image. Figure 1 shows an example of the combined FITS image of A2670 from multiple exposures. The characteristic R radius of the cluster is drawn as a dashed circle centered at the A2670’s center coordinate given in Table 1. The FoV of MOSAICII which was used by S12 is also overlaid as a thin solid square, for comparison. MOSAICII is a predecessor of the DECam which was mounted on the Blanco telescope.
| Cluster ID | R.A. | Decl. | Redshift | Filter | Date | MW Extinction | |
@@
We make use of pipeline-reduced science, weight, and data quality map (DQM) FITS image data of the clusters for both single-exposure and combined mosaics from the NOIRLab Astro Data Archive333http://astroarchive.noirlab.edu. The pipeline reduces raw DECam data to single-exposure mosaics by removing instrument signature, applying world coordinate system (WCS) and photometric calibration, and re-projecting to a common grid. The combined mosaics are then stacked and/or tiled in a region of sky derived from multiple, spatially overlapping single-exposures in the same filter band. Detailed information about the reduction process can be found in Shaw et al. (2015). Basic information for the observed clusters and their FITS mosaics that we used is summarized in Table 1. Multiple standard-star fields were also observed each observing night and used for the flux calibration. Each field was observed at least twice with a large enough time intervals to measure atmospheric extinction at different airmasses (refer to the website1 for more information).
The detection of faint merger features, such as tidal tails and shells, is depth-dependent. To measure the depth of our photometric data, we added mock sources with various magnitudes at random locations in the images using a 2D Gaussian model444https://docs.astropy.org/en/stable/api/astropy.modeling.functional_models.Gaussian2D.html. The 2D Gaussian models are commonly used to model the point spread function (PSF) of astronomical images. Those are also relatively easy to generate and control the parameters especially for measuring depths of data from multiple sources. We then determined the completeness as a function of limiting magnitude by counting the number of injected sources detected by SExtractor(see Figure 2). We also downloaded and measured depths of two overlapping public surveys: the Sloan Digital Sky Survey (SDSS; Ahumada et al., 2020) and the DECam Legacy Survey (DECaLS; Dey et al., 2019). The magnitude where the detection rate drops below 90% () was around 24 mag in - and -band of our combined DECam mosaics similar to that of DECaLS. The of SDSS was around 22 mag in the same bands. The fainter magnitude limit where the detection rate drops below 1% (), on the other hand, was around 27 mag in our data in - and -band which was 1 mag (1.5 mag in the case of Figure 2) deeper than that of DECaLS. The of SDSS in - and -band was around 23 mag. We share the result for all sample clusters in three bands on the website1. 1- surface brightness limits555https://www.gnu.org/software/gnuastro/manual/html_node/Surface-brightness-limit-of-image.html, using the median of the standard deviation calculated in moving meshes in the band, are tabulated in the website1. The values are mag arcsec for the data we use in this study.
@@ 2.2 Spectroscopy
Figure 3 shows a histogram of the radial velocities from the spectroscopic surveys and the fitted 1D Gaussian curves. The A3558 mosaic also includes a second cluster, Abell A3562. We therefore split the spectroscopic catalog of the A3558 field into two subgroups ‘a3558’ and ‘a3562’ (see Section 6 and the result for A3558 in Section 7), and fit a Gaussian curve separately (see the bottom left panel in Figure 3). We derived the size and mass of each cluster using the fitted following the method outlined by S12 (see their Equation 4).
@@ 4 Identification of merging and post-merging galaxies
Accurate estimation and subtraction of background levels is crucial in the process of galaxy stamp generation because the faint tidal features around galaxies can easily be erased by either over- or under-subtraction of the background. In this work, we use the background subtraction of Galapagos-2, which best eliminates background flux while preserving faint tidal features in the galaxy images (see Appendix B for a comparison of several background fitting techniques). Galapagos-2 is an improved version of Galapagos (Häußler et al., 2013), an IDL777Interactive Data Language code based on GALFIT (Peng et al., 2002). Galapagos-2 incorporates multi-band capabilities and allows for precise measurement of the sky background by performing a multi-source 2D fitting based on the SExtractor output catalog. We also measured the background RMS in band images and computed surface brightness limits5. The results are online1. For our sample of galaxy clusters, the surface brightness limit, , was in the range 28.0–28.7 mag arcsec with a moving mesh size of 1000x1000 pixel.
@@ 7.1 Relation with the dynamical state of clusters
Figure 6 displays and within the R of each cluster as a functions of their dynamical status indicators. For each indicator, Pearson correlation coefficients () and -values are computed and shown at the bottom of each panel. Positive correlations approach 1, while negative correlations approach 1. Results are considered statistically significant with -values below 0.05. Note that the ‘relaxed’ clusters generally exhibit lower indicator values (positioned towards the left), except for concentration, , where they have higher values (top left panel; see Yuan et al., 2022, Figure 5). We observed a marginal trend suggesting higher fractions of merging galaxies in ‘relaxed’ clusters. This trend correlates positively with concentration and negatively with the others. However, the correlation remains weak (-value 0.11–0.54). No such trend was found for PM-type fractions.
@@ 7.2 Substructures
A2670: We adopt the model with the highest BIC score having 5 components and set the orange component as the main cluster. The cluster has 4%1.5 higher than the average 3%2.8 (cluster to cluster). Piraino-Cerda et al. (in press) performed the mclust analysis with the spectroscopic data from the SDSS around A2670 in a wider range of 5 and find six components. DS+ found six substructures.
A3558: Maximum number of components, nine, were found by the mclust in the field of A3558. Still, we adopt a model with six components because the second and fifth components, colored in green and brown in Figure 8, matched well with the X-ray peaks. The two X-ray peaks are the centers of two galaxy clusters, A3558 and A3562, in the Shapley Supercluster (Raychaudhury, 1990)121212http://www.atlasoftheuniverse.com/superc/shapley.html (see ‘X’ marks in the top-left panel of Figure 8). We fit the Gaussian curves for line-of-sight velocities of the two substructures separately (see Figure 3) and draw their R in the upper-left panel of Figure 8. We name the two clusters ‘a3558’ and ‘a3562’ to avoid confusion with the DECam mosaic ID ‘A3558’, which contains two clusters, and consider both the green and brown components as the main clusters. The ‘a3558’ is also known as ‘Shapley 8’12 located at the core of the Shapley Supercluster. Four other components found by mclust are defined as substructures. In the case of DS+, as mentioned in Section 6, we find substructures in ‘a3558’ and ‘a3562’ separately, and combine the result. ‘a3558’ and ‘a3562’ resulted in five substructures each.
@@ 8 Discussion and Conclusion
We visually inspected a volume-limited sample of spectroscopically-confirmed member galaxies in six Abell clusters in the redshift range 0.046 0.076 on the wide ( 3) and deep ( mag arcsec) stacked DECam images, and classified galaxies according to the indication of ongoing- (M) and post- (PM) merging features. On average: (1) galaxies that are currently undergoing a merger occur with a frequency of 4% with a decreasing frequency from the outside to the inside of the cluster; and (2) galaxies with post-merger features are found 7% preferentially inside the escape velocity region on the PPSD and substructures. This is consistent with a scenario where galaxies undergo pre-processing outside of the galaxy clusters and accrete onto the galaxy clusters, retaining their post-merger features.
We obtain a similar ongoing merger rate to S12 and O18, while McIntosh et al. (2008) reported about a dex lower rate, 0.2%, at a dark matter halo mass range greater than from the SDSS. Shallower images are almost certainly the origin of the main difference, as Yi et al. (2013) and Ji et al. (2014) clearly demonstrated. For PM galaxies, our observed rate of 6%2 (the mean and standard deviation of the PM fraction in each cluster) is approximately 15% point lower than that reported by O18 and 20% point lower than S12. S12 analyzed a sample within a red-sequence strip, using imaging data of similar depth to ours ( mag arcsec). O18 included all Hubble types brighter than with data approximately one magnitude shallower than ours. Consistent with S12 and O18, we find higher PM rates compared to M rates. This suggest one or both of the following: 1) PM features persist for longer timescales than the total duration of the merging process itself. 2) PM galaxies migrate into the cluster after the merger has occurred outside of it.
The ratio of M- to PM-type galaxies, M/PM, in this study goes from 0.58 (34/58; voted by at least two out of four) to 0.8 (16/20; voted by at least three out of four). M/PM from the S12 and O18 studies are 0.1–0.2. N-body simulations of equal-mass gas-rich disc galaxy mergers show a merging time scale of 0.5–3 Gyr between the first pass and the final merger of the nuclei (see Table 3 in Lotz et al., 2008), while Yi et al. (2013) and Ji et al. (2014) scrutinized the time scale for galaxies retaining post-merger features, reporting 4 Gyr from the imaging data with a depth of 28 mag arcsec. The range of M/PM from the two simulation studies is 0.125–0.75, which matches the observational results.
@@
Several studies highlight the impact of cluster dynamics on SF and active galactic nuclei (AGN) activity. Ferrari et al. (2006) observed enhanced SF in a colliding cluster, while Hwang & Lee (2009) detected excess SF and AGN activity within substructures of a merging binary cluster. Additionally, Sobral et al. (2015) and Stroe et al. (2015) linked SF enhancement in local merging clusters to shocks associated with the merging process. Contrasting mergers with relaxed clusters, Stroe & Sobral (2021) found an even distribution of SF galaxies within 3 Mpc of merging cluster centers but noted their rarity in relaxed clusters. Interestingly, they observed AGN primarily in relaxed cluster outskirts (1.5–3 Mpc). However, Wittman et al. (2024) found no correlation between star-forming/AGN fractions and time since pericenter. As suggested by Stroe et al. (2015), the shock fronts in merging clusters may directly promote SF by facilitating molecular gas collapse or by fueling black holes. To investigate whether cluster mergers promote galaxy mergers, we computed Pearson correlation coefficients and corresponding -values between the fraction of ongoing/post mergers within R and the six dynamical state parameters from Yuan et al. (2022). Overall, we observed weak correlations (-values 0.11–54) between the and all six parameters, with slightly higher fractions of ongoing mergers in ‘relaxed’ clusters. However, fractions of post-mergers showed mixed correlations. Our sample size may limit statistically robust conclusions. Notably, we lack ‘relaxed’ cluster samples with lower (negative) morphology index () values. Nonetheless, the consistent correlations across all six parameter spaces warrant further investigation.

### v2

@@ Abstract
We study the incidence and spatial distribution of galaxies that are currently undergoing gravitational merging (M) or that have signs of a post merger (PM) in six galaxy clusters (A754, A2399, A2670, A3558, A3562, and A3716) within the redshift range, 0.05 0.08. To this aim, we obtained Dark Energy Camera (DECam) mosaics in , , and -bands covering up to of the clusters, reaching 28 mag/arcsec2 surface brightness limits. We visually inspect color-composite images of volume-limited () cluster-member galaxies to identify whether galaxies are of M or PM types. We find 4% M-type and 7% PM-type galaxies in the galaxy clusters studied. By adding spectroscopic data and studying the projected phase space diagram (PPSD) of the projected clustocentric radius and the line-of-sight velocity, we find that PM-type galaxies are more virialized than M-type galaxies, having 1–5% point higher fraction within the escape-velocity region, while the fraction of M-type was 10% point higher than PM-type in the intermediate environment. Similarly, on a substructure analysis, M types were found in the outskirt groups, while PM types populated groups in ubiquitous regions of the PPSD. Adopting literature-derived dynamical state indicator values, we observed a higher abundance of M types in dynamically relaxed clusters. This finding suggests that galaxies displaying post-merging features within clusters likely merged in low-velocity environments, including cluster outskirts and dynamically relaxed clusters.
2024 April 30
@@ 1 Introduction
The importance of mergers in driving the evolution of galaxies within clusters, where effects such as ram-pressure stripping (Gunn & Gott, 1972) and tidal interactions (Jones et al., 2000; Gnedin, 2003) begin to play an increasingly important role, is a matter of discussion. Although clusters represent particularly dense environments, the incidence of mergers among their members is not expected to be high due to high relative velocities. Analytically, the energy input and dynamical friction of a collision scale with the inverse square of the relative velocity of the interacting galaxies (Binney & Tremaine, 1987), such that fast encounters, which become preponderant in denser cluster environments, are less likely to result in a merger than slow encounters, which are more common in the field (Makino & Hut, 1997; Mihos, 2003). Studies such as Sheen et al. (2012) (S12 hereafter) and Oh et al. (2018) (O18 hereafter) report, indeed, that only around four percent of the early-type galaxies (ETGs) found in cluster environments appear to be actively undergoing major mergers. Oh et al. (2018) report only a marginal change in the fraction of ongoing mergers, which increases from 4% to 7.7% when going from 2 to 2 in clustocentric distance, concluding that merger processes are unlikely to significantly impact galaxy evolution within the virialized region of a cluster. McIntosh et al. (2008) found also negative correlation of the major merging frequency of massive satellite galaxies with the host halo mass. Their sample of SDSS satellite galaxies with a stellar mass M⊙ dropped their major merging frequency from 3% in halos with their mass around to almost 0% in cluster-scale halos with .
@@
Given the apparently negligible merger rates in cluster environments, the high proportion of cluster galaxies that exhibit post-merging features is therefore unexpected. Deep optical surveys of galaxies in dense cluster environments show a high fraction of post-merging features (25% of galaxies brighter than 20 in the band in the red sequences in S12; 20% of galaxies with similar brightnesses in O18), with abundances comparable to that of the field (see e.g. van Dokkum, 2005; Kado-Fong et al., 2018; Bílek et al., 2020; Sola et al., 2022; Trujillo et al., 2021; Martin et al., 2022; Valenzuela & Remus, 2022). S12 attribute this result to migration of galaxy populations that have undergone mergers prior to entering the cluster (pre-processing; Fujita, 2004). The work of Yi et al. (2013) and Ji et al. (2014) lend support to this argument, showing that, merger features continue to be detectable for about 4 Gyr at a limiting surface brightness of 28 mag arcsec-2 in their hydrodynamic simulations of major mergers between Sa- and Sb-type galaxies. Evidence of pre-processing is also seen at higher redshifts. For instance, van Dokkum et al. (1999) find that merging galaxies are found preferentially in the outskirts of galaxy clusters at , and Olave-Rojas et al. (2018) find high quenching efficiencies in galaxies at large distances from the cluster centers at .
@@
Besides, clusters are consistently accreting groups. Das et al. (2023) found a higher level of star formation (SF) and bluer colors in low-mass SDSS galaxy pairs in filaments and sheets around galaxy clusters. Kleiner et al. (2014) found more asymmetric galaxies in the inner region of A1664 where a merging group most likely passed the pericenter, and they argued that a merger between halos of a group and a cluster enhanced galaxy-galaxy interactions (see also Vijayaraghavan & Ricker, 2013). Mihos (2003) also stressed that the infalling galaxy groups through the “cosmic web” would allow galaxies to interact with each other strongly by slow encounters. Łokas (2023) also found galaxy-galaxy interactions in a galaxy group prior to cluster infall in the IllustridTNG100 simulation. Analyses of substructures in/outside of clusters seem to be necessary to study galaxy-galaxy interactions in merging clusters (i.e., Olave-Rojas et al., 2023).
@@
In this paper, we study the abundances of gravitationally interacting or post-merger galaxies using deep (28 mag arcsec-2)5 and wide (3 deg2; 3) photometric data sets (in , , and ) of six Abell clusters: A754, A2399, A2670, A3558, A3562, and A3716 in a redshift range, 0.05 0.08 collected using the Dark Energy Camera (DECam; DePoy et al., 2008). In combination with public spectroscopic data, we study the prevalence of these features in substructures and in regions of the projected phase-space diagram. This is done to understand the evolution of galaxies over the course of their accretion onto the clusters.
@@ 2.1 Photometric measurements
Optical mosaics in , , and bands of the field of views (FoVs) around seven Abell galaxy clusters were taken using the Dark Energy Camera (DECam; e.g., Flaugher et al., 2015) installed on the Blanco 4m Telescope at CTIO from April 10th to 11th in 2013 (Proposal ID: 2013A-0612, PI: Y. Sheen) and August 19th to 22nd in 2014 (Proposal ID: 2014B-0608, PI: Y. Jaffé), as presented in Table 1 (for A3574 and A3659, visit the website111https://data.kasi.re.kr/vo/DECam_catalogs/). DECam features a 3 deg2 FoV consisting of sixty-two 2k 4k pixel2 CCDs, with a total of 520 million pixels for each FITS222Flexible Image Transport System (Wells et al., 1981; Hanisch et al., 2001) image. Figure 1 shows an example of the combined FITS image of A2670 from multiple exposures. The characteristic R200 radius of the cluster is drawn as a dashed circle centered at the A2670’s center coordinate given in Table 1. The FoV of MOSAICII which was used by S12 is also overlaid as a thin solid square, for comparison. MOSAICII is a predecessor of the DECam which was mounted on the Blanco telescope.
Cluster ID | R.A.a | Decl.a | Redshifta | Filter | b | Dateb | MW Extinctionc |
@@
We make use of pipeline-reduced science, weight, and data quality map (DQM) FITS image data of the clusters for both single-exposure and combined mosaics from the NOIRLab Astro Data Archive333http://astroarchive.noirlab.edu. The pipeline reduces raw DECam data to single-exposure mosaics by removing instrument signature, applying world coordinate system (WCS) and photometric calibration, and re-projecting to a common grid. The combined mosaics are then stacked and/or tiled in a region of sky derived from multiple, spatially overlapping single-exposures in the same filter band. Detailed information about the reduction process can be found in Shaw & Swaters (2015). Basic information for the observed clusters and their FITS mosaics that we used is summarized in Table 1. Multiple standard-star fields were also observed each observing night and used for the flux calibration. Each field was observed at least twice with a large enough time intervals to measure atmospheric extinction at different airmasses (refer to the website1 for more information).
The detection of faint merger features, such as tidal tails and shells, is depth-dependent. To measure the depth of our photometric data, we added mock sources with various magnitudes at random locations in the images using a 2D Gaussian model444https://docs.astropy.org/en/stable/api/astropy.modeling.functional_models.Gaussian2D.html. The 2D Gaussian models are commonly used to model the point spread function (PSF) of astronomical images. Those are also relatively easy to generate and control the parameters especially for measuring depths of data from multiple sources. We then determined the completeness as a function of limiting magnitude by counting the number of injected sources detected by SExtractor(see Figure 2). We also downloaded and measured depths of two overlapping public surveys: the Sloan Digital Sky Survey (SDSS; Ahumada et al., 2020) and the DECam Legacy Survey (DECaLS; Dey et al., 2019). The magnitude where the detection rate drops below 90% () was around 24 mag in - and -band of our combined DECam mosaics similar to that of DECaLS. The of SDSS was around 22 mag in the same bands. The fainter magnitude limit where the detection rate drops below 1% (), on the other hand, was around 27 mag in our data in - and -band which was 1 mag (1.5 mag in the case of Figure 2) deeper than that of DECaLS. The of SDSS in - and -band was around 23 mag. We share the result for all sample clusters in three bands on the website1. 1- surface brightness limits555https://www.gnu.org/software/gnuastro/manual/html_node/Surface-brightness-limit-of-image.html, using the median of the standard deviation calculated in moving meshes in the band, are tabulated in the website1. The values are mag arcsec-2 for the data we use in this study.
@@ 2.2 Spectroscopy
Figure 3 shows a histogram of the radial velocities from the spectroscopic surveys and the fitted 1D Gaussian curves. The A3558 mosaic also includes a second cluster, A3562. We therefore split the spectroscopic catalog of the A3558 field into two subgroups ‘a3558’ and ‘a3562’ (see Section 6 and the result for A3558 in Section 7), and fit a Gaussian curve separately (see the bottom left panel in Figure 3). We derived the size and mass of each cluster using the fitted following the method outlined by S12 (see their Equation 4).
@@ 4 Identification of merging and post-merging galaxies
Accurate estimation and subtraction of background levels is crucial in the process of galaxy stamp generation because the faint tidal features around galaxies can easily be erased by either over- or under-subtraction of the background. In this work, we use the background subtraction of Galapagos-2, which best eliminates background flux while preserving faint tidal features in the galaxy images (see Appendix B for a comparison of several background fitting techniques). Galapagos-2 is an improved version of Galapagos (Häußler et al., 2013), an IDL777Interactive Data Language code based on GALFIT (Peng et al., 2002). Galapagos-2 incorporates multi-band capabilities and allows for precise measurement of the sky background by performing a multi-source 2D fitting based on the SExtractor output catalog. We also measured the background RMS in band images and computed surface brightness limits5. The results are online1. For our sample of galaxy clusters, the surface brightness limit, , was in the range 28.0–28.7 mag arcsec-2 with a moving mesh size of 1000x1000 pixel2.
@@ 7.1 Relation with the dynamical state of clusters
Figure 6 displays and within the R200 of each cluster as a functions of their dynamical status indicators. For each indicator, Pearson correlation coefficients () and -values are computed and shown at the bottom of each panel. Positive correlations approach 1, while negative correlations approach 1. Results are considered statistically significant with -values below 0.05. Note that the ‘relaxed’ clusters generally exhibit lower indicator values (positioned towards the left), except for concentration, , where they have higher values (top left panel; see Yuan et al., 2022, Figure 5). We observed a marginal trend suggesting higher fractions of merging galaxies in ‘relaxed’ clusters. This trend correlates positively with concentration and negatively with the others. However, the correlation remains weak (-value 0.11–0.54). No such trend was found for PM-type fractions.
@@ 7.2 Substructures
A2670: We adopt the model with the highest BIC score having 5 components and set the orange component as the main cluster. The cluster has 4%1.5 higher than the average 3%2.8 (cluster to cluster). Piraino-Cerda et al. (2023) performed the mclust analysis with the spectroscopic data from the SDSS around A2670 in a wider range of 5 and find six components. DS+ found six substructures.
A3558: Maximum number of components, nine, were found by the mclust in the field of A3558. Still, we adopt a model with six components because the second and fifth components, colored in green and brown in Figure 8, matched well with the X-ray peaks. The two X-ray peaks are the centers of two galaxy clusters, A3558 and A3562, in the Shapley Supercluster (Raychaudhury, 1990)121212http://www.atlasoftheuniverse.com/superc/shapley.html (see ‘X’ marks in the top-left panel of Figure 8). We fit the Gaussian curves for line-of-sight velocities of the two substructures separately (see Figure 3) and draw their R200 in the upper-left panel of Figure 8. We name the two clusters ‘a3558’ and ‘a3562’ to avoid confusion with the DECam mosaic ID ‘A3558’, which contains two clusters, and consider both the green and brown components as the main clusters. The ‘a3558’ is also known as ‘Shapley 8’12 located at the core of the Shapley Supercluster. Four other components found by mclust are defined as substructures. In the case of DS+, as mentioned in Section 6, we find substructures in ‘a3558’ and ‘a3562’ separately, and combine the result. ‘a3558’ and ‘a3562’ resulted in five substructures each.
@@ 8 Discussion and Conclusion
We visually inspected a volume-limited sample of spectroscopically-confirmed member galaxies in six Abell clusters in the redshift range 0.046 0.076 on the wide ( 3) and deep ( mag arcsec-2) stacked DECam images, and classified galaxies according to the indication of ongoing- (M) and post- (PM) merging features. On average: (1) galaxies that are currently undergoing a merger occur with a frequency of 4% with a decreasing frequency from the outside to the inside of the cluster; and (2) galaxies with post-merger features are found 7% preferentially inside the escape velocity region on the PPSD and substructures. This is consistent with a scenario where galaxies undergo pre-processing outside of the galaxy clusters and accrete onto the galaxy clusters, retaining their post-merger features.
We obtain a similar ongoing merger rate to S12 and O18, while McIntosh et al. (2008) reported about a dex lower rate, 0.2%, at a dark matter halo mass range greater than from the SDSS. Shallower images are almost certainly the origin of the main difference, as Yi et al. (2013) and Ji et al. (2014) clearly demonstrated. For PM galaxies, our observed rate of 6%2 (the mean and standard deviation of the PM fraction in each cluster) is approximately 15% point lower than that reported by O18 and 20% point lower than S12. S12 analyzed a sample within a red-sequence strip, using imaging data of similar depth to ours ( mag arcsec-2). O18 included all Hubble types brighter than with data approximately one magnitude shallower than ours. Consistent with S12 and O18, we find higher PM rates compared to M rates. This suggest one or both of the following: 1) PM features persist for longer timescales than the total duration of the merging process itself. 2) PM galaxies migrate into the cluster after the merger has occurred outside of it.
The ratio of M- to PM-type galaxies, M/PM, in this study goes from 0.58 (34/58; voted by at least two out of four) to 0.8 (16/20; voted by at least three out of four). M/PM from the S12 and O18 studies are 0.1–0.2. N-body simulations of equal-mass gas-rich disc galaxy mergers show a merging time scale of 0.5–3 Gyr between the first pass and the final merger of the nuclei (see Table 3 in Lotz et al., 2008), while Yi et al. (2013) and Ji et al. (2014) scrutinized the time scale for galaxies retaining post-merger features, reporting 4 Gyr from the imaging data with a depth of 28 mag arcsec-2. The range of M/PM from the two simulation studies is 0.125–0.75, which matches the observational results.
@@
Several studies highlight the impact of cluster dynamics on SF and active galactic nuclei (AGN) activity. Ferrari et al. (2006) observed enhanced SF in a colliding cluster, while Hwang & Lee (2009) detected excess SF and AGN activity within substructures of a merging binary cluster. Additionally, Sobral et al. (2015) and Stroe et al. (2015) linked SF enhancement in local merging clusters to shocks associated with the merging process. Contrasting mergers with relaxed clusters, Stroe & Sobral (2021) found an even distribution of SF galaxies within 3 Mpc of merging cluster centers but noted their rarity in relaxed clusters. Interestingly, they observed AGN primarily in relaxed cluster outskirts (1.5–3 Mpc; see also Koulouridis et al., 2018, 2024; Koulouridis & Bartalucci, 2019). However, Wittman et al. (2024) found no correlation between star-forming/AGN fractions and time since pericenter. As suggested by Stroe et al. (2015), the shock fronts in merging clusters may directly promote SF by facilitating molecular gas collapse or by fueling black holes. To investigate whether cluster mergers promote galaxy mergers, we computed Pearson correlation coefficients and corresponding -values between the fraction of ongoing/post mergers within R200 and the six dynamical state parameters from Yuan et al. (2022). Overall, we observed weak correlations (-values 0.11–54) between the and all six parameters, with slightly higher fractions of ongoing mergers in ‘relaxed’ clusters. However, fractions of post-mergers showed mixed correlations. Our sample size may limit statistically robust conclusions. Notably, we lack ‘relaxed’ cluster samples with lower (negative) morphology index () values. Nonetheless, the consistent correlations across all six parameter spaces warrant further investigation.